Manual for Theses and Dissertations formatted

with ua-thesis.cls with the draft Option

by

The Graduate College

A Dissertation Submitted to the Faculty of the

Graduate Interdisciplinary Program

in Applied Mathematics

In Partial Fulfillment of the Requirements

For the Degree of

Doctor of Philosophy

In the Graduate College

The University of Arizona

1 9 9 6

Acknowledgments

The contents of this example dissertation has been entirely written by the Graduate College at the University of Arizona.

Table of Contents

List of Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

List of Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Abstract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Chapter 1. Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . 8

Chapter 2. University Microfilms Incorporated (UMI) . . . . . . 9

Chapter 3. General Format Requirements . . . . . . . . . . . . . . 10 3.1. Margins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2. Corrections on Pages . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.3. Page numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.4. Paper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.5. Photocopy Quality . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.6. Printers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.7. Type Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

3.7.1. Type Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.7.2. Typewritten Papers . . . . . . . . . . . . . . . . . . . . . . . . 11

Chapter 4. Order of Sections . . . . . . . . . . . . . . . . . . . . . . . 12

4.1. Sample Title Page . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.1.1. Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.1.2. Copyrighting the Thesis or Dissertation . . . . . . . . . . . . . 17

4.2. Statement by Author . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 4.3. Notes for Table of Contents . . . . . . . . . . . . . . . . . . . . . . . 17 4.4. Notes for List of Illustrations / List of Tables . . . . . . . . . . . . . 22 4.5. Abstract and Special Abstract Compared . . . . . . . . . . . . . . . . 23

4.5.1. Abstract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 4.5.2. Special Abstract . . . . . . . . . . . . . . . . . . . . . . . . . 23

Appendix A. Inclusion of Published Papers or Manuscripts for Publication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 A.1. Body of Paper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 A.2. Appendices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Appendix B. Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Appendix C. Human/Animal Subjects Approval . . . . . . . . . . . . 30 Appendix D. Illustrations, Tables, Graphs . . . . . . . . . . . . . . . 31 Appendix E. Oversized Materials . . . . . . . . . . . . . . . . . . . . . 32 Appendix F. Photographs . . . . . . . . . . . . . . . . . . . . . . . . . . 33 References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

List of Figures

Figure 4.1. Example Title Page . . . . . . . . . . . . . . . . . . . . . . . . 16 Figure 4.2. Non-copyrighted thesis . . . . . . . . . . . . . . . . . . . . . . . 18 Figure 4.3. Copyrighted thesis . . . . . . . . . . . . . . . . . . . . . . . . . 19 Figure 4.4. Non-copyrighted dissertation . . . . . . . . . . . . . . . . . . . 20 Figure 4.5. Copyrighted dissertation . . . . . . . . . . . . . . . . . . . . . . 21 Figure 4.6. Example 1 for Table of Contents . . . . . . . . . . . . . . . . . 22 Figure 4.7. Example 2 for Table of Contents . . . . . . . . . . . . . . . . . 22 Figure 4.8. Sample List of Illustrations . . . . . . . . . . . . . . . . . . . . 23 Figure 4.9. Sample List of Tables . . . . . . . . . . . . . . . . . . . . . . . 23 Figure 4.10. Sample Abstract . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Figure 4.11. Sample Special Abstract . . . . . . . . . . . . . . . . . . . . . . 26

List of Tables

Table 3.1. Required Paper . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

Abstract

This example dissertation contains the original text of the “Manual for Theses and Dissertations”, written by the Graduate College at the University of Arizona. It has been obtained via the internet at

http://grad.admin.arizona.edu/degreecert/ThesisManual/manual.htm

on May 10, 1996. The page was last updated November 9, 1995. No guarantee is made that this information is current, and students should check with the Graduate College before submitting a dissertation or thesis.

Chapter 1

Introduction

Use this manual as a guide for setting up the physical format of your thesis, dissertation or document. Your thesis will represent you, your department, and The University of Arizona in the international scholarly community. Your work is important and worthy of professional presentation. This manual lists Graduate College requirements for the mechanical aspects of meeting these high standards.

In this manual the word thesis, includes documents and dissertations. If format requirements for the document or dissertation vary from those for the thesis, specific requirements for each type of paper will be listed.

Two final copies of the thesis must be submitted; both must meet all specifications of this manual. The two final copies should be submitted unbound in a box to the Graduate College Degree Certification Office.

Chapter 2

University Microfilms Incorporated (UMI)

Your thesis will be published by University Microfilms Incorporated, Ann Arbor, Michigan. Upon certification by your major professor, your examining committee, and the Graduate College, a copy of the thesis and a Special Abstract are forwarded to UMI. The manuscript is cataloged and microfilmed, the microfilm negative is inspected and put in vault storage. Paper copies of your work will be produced on demand by UMI. Catalog information is sent to the Library of Congress for production and distribution of catalog cards for libraries. The original copy of the thesis is returned to The University of Arizona Library. The Special Abstract is printed in Microfilm Abstracts and distributed to leading libraries in the United States and abroad and to a selected list ofjournals and abstracting services.

Publication by UMI does not preclude publication by other means later. You are urged to submit your work for publication in a scholarly or professional journal. Suitable acknowledgment must indicate that the publication is a thesis, dissertation, or document, or portion thereof, which was submitted in partial fulfillment of the requirements for a degree at the University of Arizona.

You must complete a UMI publication agreement, available through the Degree Certification Office.

Chapter 3

General Format Requirements

3.1 Margins

Text, illustrations (figures) or tables must not appear outside the specified margins. Specific margin requirements are listed in ORDER OF SECTIONS under each category. Page numbers are the only item which may appear outside the margin requirements.

3.2 Corrections on Pages

Do not use correction fluid or correction tape. These materials flake off in handling and storage, exposing the original errors.

3.3 Page numbers

The title page is page 1 of the thesis. All pages which follow are numbered in a single sequence with arabic numerals. Page numbers must be placed at least 1 ” below the top of the sheet, and 1” from the right edge. The numbers must be at least 1/4” above the first line of text. You may omit the printed page number on the title page; all other pages must have printed page numbers. Do not use page headers. Do not use the phrase, Page xx; just the numeral.

3.4 Paper

See Table 3.1.

1.  Xerox Image Series Smooth Paper
2.  Xerox Image Elite Paper 25% Cotton
3.  Xerox Series 10
4.  100% Cotton - 20 lb weight
5.  Cranes Crest Bond - 100% Cotton Acid Free

Table 3.1. Required Paper

3.5 Photocopy Quality

Photocopies must meet all requirements for margins, readability, and type of paper. This includes all photocopied documents, tables, illustrations and appendix pages.

3.6 Printers

Laser printing or other letter quality printing is required. Impact, or daisy wheel printing is generally acceptable. 24-pin dot matrix near letter quality and draft quality printing are not acceptable.

3.7 Type Fonts

Standard serif typefaces such as Courier and Times Roman reproduce and microfilm well. Do not use modern Sans Serif types, which read well in the original but do not reduce well for microfilming. Ornamental styles such as Script and Old English may not be used. Limit the use of italic styles to standard uses in bibliographic citations and foreign words. Boldface should be restricted to very small segments of the text and to infrequent occurrences.

3.7.1 Type Size

Use 12-point or 14-point for proportional fonts; 10 cpi or 12 cpi for non-proportional fonts. A proportional font allows proportional spacing - a feature that gives a printed page a more pleasing appearance by allowing for different widths of characters. The letter w, for example, is wider than the letter i. Normally, when these letters are printed, both are given the same amount of space; the result can be gaps that are visually distracting. With proportional printing, the letter w is given more space than the letter i, creating a more aesthetic and professional-looking line of text.

3.7.2 Typewritten Papers

Papers prepared on good quality electric typewriters are acceptable. All margin, paper quality, and typographic requirements apply. Type size should be Pica (10 cpi) or Elite (12 cpi)

Chapter 4

Order of Sections

Components of your thesis must be in the following order, formatted as specified:

1. Title Page • Required

• Margins:

– Top 2.5”

– Bottom 1.5”

– Left 1.5”

– Right 1”

• Spacing: Follow sample

2. Final Examining Committee Approval Form

• Required for dissertations and music documents, not for theses. The approval form for the thesis is included in the Statement by Author (see item 3 below).

• Note: Before the final oral defense the student obtains the Approval Pages from the Degree Certification Office. Original signatures are required on both final copies.

3. Statement by Author • Required

• Margins:

– Top 2.5”

– Bottom 1”

– Left 1.5”

– Right 1”

• Spacing: Single

• Note: Follow examples. Original signatures are required on both final

copies.

4. Acknowledgements • Optional

• Margins: Same as Body of Paper

• Spacing: Maybe single spaced

• Note: One page maximum

5. Dedication • Optional

• Margins: Same as Body of Paper

• Spacing: Must be double spaced

• Note: One page maximum

6. Table of Contents • Required

• Margins: Same as Body of Paper

• Note: See ”Notes for Table of Contents” for notes on format.

7. List of Illustrations/List of Tables • Required if document contains illustrations, figures or tables.

• Margins: Same as Body of Paper

• Note: Formatted like Table of Contents; see ”Notes for List of Illustrations

/ List of Tables” for more information.

8. Abstract • Required

• Margins:

– Top 1.5”

– Bottom 1”

– Left 1.5”

– Right 1”

• Spacing: Double spaced

• Note: A Special Abstract for UMI is also required. The text remains the same for both versions, but formatting requirements differ. See ”Abstract and Special Abstract Compared”.

9. Body of Paper • Required

• Margins:

– Top 1.5”

– Bottom 1”

– Left 1.5”

– Right 1”

• Spacing: Double, except for long quotations, footnotes, table and illustration captions

• Note: Begin each major section on a new page. Margin requirements apply to every page of the thesis unless otherwise specified in this manual. See APPENDIX A, INCLUSION OF PUBLISHED PAPERS OR MANUSCRIPTS FOR PUBLICATION, if your department allows this option.

10. Appendices • Optional

• Margins: Same as Body of Paper

• Spacing: Depends on nature of Appendix material

• Note: Each Appendix must begin on a new page.

11. References • Required if citations are used

• Margins: Same as Body of Paper

• Spacing: Citations single spaced; double space between citations

• Note: Use your department’s preferred citation style; consult with your advisor if more than one style is acceptable. Title this section REFER-ENCES or WORKS CITED. Do not use the word, Bibliography.

4.1 Sample Title Page

Margins:

• Top 2.5”

• Bottom 1.5”

• Left 1.5”

• Right 1”

• The title page is centered within the left and right margins. Title in capital letters:

               SEBASTIEN CASTELLIO, APOSTLE OF TOLERANCE
                       IN THE SIXTEENTH CENTURY

Use your full name, spelled out:

                                 by
                         Jane Allison Smith

This rule (solid line) is 2” long and is placed approximately 5” below the top of the page. Copyright statement, if used, is placed directly below the rule:

                       _____________________
                Copyright (c) Jane Allison Smith 1992

Follow the capitalization and spacing of the lines in the bottom section:

             A Thesis Submitted to the Faculty of the
                 DEPARTMENT OF ROMANCE LANGUAGES
           In Partial Fulfillment of the Requirements
                       For the Degree of
                        MASTER OF ARTS
                   WITH A MAJOR IN SPANISH
                   In the Graduate College
                  THE UNIVERSITY OF ARIZONA
                          1 9 9 2

4.1.1 Notes

This is a sample page for a THESIS. For a doctoral degree substitute DISSERTA-TION for THESIS; for a musical arts degree, substitute DOCUMENT for THESIS. Substitute DOCTOR OF PHILOSOPHY, DOCTOR OF MUSICAL ARTS, or other title as appropriate.

The statement, WITH A MAJOR IN..., is included only if the name of the major field of study is not exactly the same as the official name of the department

Put spaces between the digits of the year: 1 (space) 9 (space) 9 (space) 2 In its entirety, see Figure 4.1

SEBASTIEN CASTELLIO, APOSTLE OF TOLERANCE
IN THE SIXTEENTH CENTURY
           by
   Jane Allison Smith
_____________________
Copyright  Jane Allison Smith 1992
A Thesis Submitted to the Faculty of the
DEPARTMENT OF ROMANCE LANGUAGES
In Partial Fulfillment of the Requirements
For the Degree of
 MASTER OF ARTS
WITH A MAJOR IN SPANISH
In the Graduate College
THE UNIVERSITY OF ARIZONA
    1 9 9 2

Figure 4.1. Example Title Page 4.1.2 Copyrighting the Thesis or Dissertation

Copyrighting of a thesis is optional. Publication by University Microfilms does not preclude publication by other methods later. If you want University Microfilms Incorporated to file, on your behalf, an application for registration of a claim of copyright on your manuscript, you must indicate this on the agreement form you complete for UMI and submit the required fee by certified check or money order. This service includes payment of the registration fee, preparation of the application, and submission of copies required by the Copyright Office.

The ownership of a copyright shall reside with the student unless otherwise stated by University policy or by terms of the research grants, fellowships, financial aid, etc. which were used to support the student’s research.

Additional information on obtaining a copyright is available from the Graduate College Degree Certification Office or the United States Copyright Office, Library of Congress, Washington, D. C. 20559.

4.2 Statement by Author

Below are four samples of the Statement by Author. Select the applicable sample: non-copyrighted thesis, copyrighted thesis, non-copyrighted dissertation, or copyrighted dissertation, and copy it. Doctor of Musical Arts students must substitute document for thesis or dissertation. For theses, substitute the full name of your advisor for the name on the sample page, Figures 4.2, 4.3, 4.4, 4.5.

Margins:

• Top 2.5”

• Bottom 1.5”

• Left 1.5”

• Right 1”

4.3 Notes for Table of Contents

The Table of Contents is required. All levels of subheadings in your manuscript must appear in the Table of Contents. The TABLE OF CONTENTS for this Manual contains two levels; your table of contents might contain more.

Margins:

• Top 1.5”

• Bottom 1”

                 STATEMENT BY AUTHOR
   This thesis has been submitted in partial fulfillment of
requirements for an advanced degree at The University of
Arizona and is deposited in the University Library to be
made available to borrowers under rules of the Library.
   Brief quotations from this thesis are allowable without
special permission, provided that accurate acknowledgment of
source is made.  Requests for permission for extended
quotation from or reproduction of this manuscript in whole
or in part may be granted by the head of the major
department or the Dean of the Graduate College when in his
or her judgment the proposed use of the material is in the
interests of scholarship.  In all other instances, however,
permission must be obtained from the author.
                     SIGNED: ________________________________
               APPROVAL BY THESIS DIRECTOR
  This thesis has been approved on the date shown below:
__________________________________________________________
         Jane M. Doe                        Date
   Professor of Chemistry

Figure 4.2. Non-copyrighted thesis

                 STATEMENT BY AUTHOR
   This thesis has been submitted in partial fulfillment of
requirements for an advanced degree at The University of
Arizona and is deposited in the University Library to be
made available to borrowers under rules of the Library.
   Brief quotations from this thesis are allowable without
special permission, provided that accurate acknowledgment of
source is made.  Requests for permission for extended
quotation from or reproduction of this manuscript in whole
or in part may be granted by the copyright holder.
                     SIGNED: ________________________________
               APPROVAL BY THESIS DIRECTOR
  This thesis has been approved on the date shown below:
__________________________________________________________
         Jane M. Doe                        Date
   Professor of Chemistry

Figure 4.3. Copyrighted thesis

                 STATEMENT BY AUTHOR
   This dissertation has been submitted in partial
fulfillment of requirements for an advanced degree at The
University of Arizona and is deposited in the University
Library to be made available to borrowers under rules of the
Library.
   Brief quotations from this dissertation are allowable
without special permission, provided that accurate
acknowledgment of source is made.  Requests for permission
for extended quotation from or reproduction of this
manuscript in whole or in part may be granted by the head of
the major department or the Dean of the Graduate College
when in his or her judgment the proposed use of the material
is in the interests of scholarship.  In all other instances,
however, permission must be obtained from the author.
                     SIGNED: ________________________________

Figure 4.4. Non-copyrighted dissertation

                 STATEMENT BY AUTHOR
   This dissertation has been submitted in partial
fulfillment of requirements for an advanced degree at The
University of Arizona and is deposited in the University
Library to be made available to borrowers under rules of the
Library.
   Brief quotations from this dissertation are allowable
without special permission, provided that accurate
acknowledgment of source is made.  Requests for permission
for extended quotation from or reproduction of this
manuscript in whole or in part may be granted by the copyright
holder.
                     SIGNED: ________________________________

Figure 4.5. Copyrighted dissertation

• Left 1.5”

• Right 1”

Format requirements include:

• The heading TABLE OF CONTENTS at the top of the first page of this section, and TABLE OF CONTENTS - Continued on each continuation page.

• Page numbers at upper right.

• Dot leaders .......................... from headings to page numbers.

• Indent each level of subheadings 4 spaces from the level above.

• Headings in the Table of Contents must exactly match the headings used in the body of the paper, and should be typographically the same (e.g., type font and style, capitalization).

• Use all capital letters for major headings. (Subheadings may be upper and lower case.)

• Each Appendix must have its own letter designation and title. Appendices are major divisions. The title appears in caps on the left margin at the same level of importance as chapter headings.

       I. LIST OF TABLES ...............................  6
      II. INTRODUCTION .................................  8
              Literature Review ........................  8
              Statistical Methods ...................... 11

Figure 4.6. Example 1 for Table of Contents

       1. LIST OF TABLES ...............................  6
       2. INTRODUCTION .................................  8
             2.1 Literature Review .....................  8
             2.2 Statistical Methods ................... 11

Figure 4.7. Example 2 for Table of Contents

Format options include:

• Chapter numbering. You may number your chapters with either arabic or roman numerals.

• Subheading numbers. If chapters are numbered, you may also number subheadings.

For example, see Figures 4.6 and 4.7.

4.4 Notes for List of Illustrations / List of Tables

These lists, which resemble the Table of Contents, are required if your thesis contains illustrations, figures, graphs or tables. Include a List of Illustrations (or List of Figures) for figures, maps and drawings. Include a List of Tables for graphs and tables. Illustrations or tables which appear in the appendices only may or may not be included with the List of Illustrations (or List of Figures) or the List of Tables.

Material in the List of Illustrations is numbered in sequence, Figure 1, Figure 2, etc. You may construct this sequence as you wish, e.g., Figure 1.1, 1.2, 2.1, 2.2.... Use LIST OF ILLUSTRATIONS as the title for the first page and LIST OF ILLUSTRATIONS - Continued for subsequent pages. You may use LIST OF FIGURES instead of LIST OF ILLUSTRATIONS if your department prefers.

Material in the List of Tables should be given its own separate sequence of numbers, Table 1, Table 2, etc. You may also construct this sequence as you wish. Use LIST OF TABLES as the title for the first page and LIST OF TABLES - Continued for subsequent pages.

                         LIST OF ILLUSTRATIONS
FIGURE 1.1, Topographic map of valley ............................. 14
FIGURE 1.2, View of valley, northwest approach .................... 20
FIGURE 2.1, Cross section of burial mound ......................... 27
FIGURE 2.2, Intact jar found near mound ........................... 28

Figure 4.8. Sample List of Illustrations

                              LIST OF TABLES
TABLE 1.1, Population estimates, 850-1100 a.d. .................... 16
TABLE 2.1, Number and type of artifacts compared with nearby sites  29

Figure 4.9. Sample List of Tables

For examples see Figures 4.8 and 4.9.

4.5 Abstract and Special Abstract Compared

4.5.1 Abstract

The Abstract is included as a page of your thesis. It is a numbered page in the thesis, appearing just before the main body of the text. The heading ABSTRACT is centered at the beginning of the first page. A sample abstract page follows (Figure 4.10). Compare it with the Special Abstract sample.

Margins:

• Top 1.5”

• Bottom 1”

• Left 1.5”

• Right 1”

4.5.2 Special Abstract

The Special Abstract contains the same text as the Abstract, but is formatted for microfilming by UMI for Dissertation Abstracts International. Treat the Special Abstract as a separate document. Include one copy of the Special Abstract and two extra copies of your title page when submitting the two final copies of your thesis.

Note the heading of the Special Abstract sample page which follows. Use the thesis title as it appears on your title page. Use your full name, and add the appropriate designation for your degree. Include your director’s name as shown (Figure 4.11).

Margins:

• Top 1.5”

• Bottom 1”

• Left 1.5”

• Right 1”

                         ABSTRACT
   This is the body of your abstract, limited to 150 words
for a thesis, and 350 words for a dissertation or document.
The word count limits apply to the regular Abstract in the
thesis and to this separate Special Abstract.  Use the same
text for both; just adjust the margins and heading.  The
abstract should summarize your work.  The UMI booklet listed
in the resources section of this Manual (here) provides some
writing tips.  The abstract for a dissertation or document
may be longer than one page; word count is more important
than page length in this section.
   This version of the Abstract, with simple heading, page
number, and 1.5" top margin, is included as part of your
thesis.

Figure 4.10. Sample Abstract

                      COMPLETE TITLE OF
            YOUR THESIS, DISSERTATION OR DOCUMENT
                   Jane Allison Doe, M.S.
               The University of Arizona, 1992
Director: John J. Jones
   This is the body of your abstract, limited to 150 words
for a thesis, and 350 words for a dissertation or document.
The word count limits apply to the regular Abstract in the
thesis and to this separate Special Abstract.  Use the same
text for both; just adjust the margins and heading.  The
abstract should summarize your work.  The UMI booklet listed
in the resources section of this Manual (here) provides some
writing tips.  The abstract for a dissertation or document
may be longer than one page; word count is more important
than page length in this section.
   This version of the Abstract, with special heading, no
page number, and 1.5" top margin, is a separate document for
UMI.  Submit one copy of the special abstract, and two extra
copies of your title page, in the box with the final copies
of your thesis.

Figure 4.11. Sample Special Abstract

Appendix A

Inclusion of Published Papers or Manuscripts for Publication

Under a policy adopted by The University of Arizona Graduate Council in January, 1992, your department may allow published and publishable papers to be included as part of your thesis. The reprints or manuscripts are treated as appendices, and the body of your thesis must include a summary of your contribution and a summary of the research. The Graduate College will accept theses in this format from any unit with an implementation policy on file with the Graduate College Degree Certification Office.

A.1 Body of Paper

The ORDER OF SECTIONS applies. In addition, the Body of the Paper must include two chapters as follows:

1. An introduction describing the unique contribution of your work to the field of study. Use the following subsections as appropriate:

(a) Explanation of the problem and its context

(b) A review of the literature

(c) Explanation of thesis format This subsection explains the relationship of the papers included and your contribution to each of the papers; where doctoral research efforts are part of a larger collaborative project, you must be able to identify one aspect of the project as your own and demonstrate an original contribution. Your role in the research and production of the published paper(s) should be clearly specified.

2. A chapter titled PRESENT STUDY which summarizes the methods, results, and conclusions of the research. The chapter should begin with a statement such as:

The methods, results, and conclusions of this study are presented in the papers appended to this thesis. The following is a summary of the most important findings in these papers.

A.2 Appendices

All mechanical requirements for Appendices listed in the ORDER OF SECTIONS apply. Your appendices will consist of:

1. A reprint of each paper as a separate appendix in the following order:

(a) a copy of the title page of the journal in which the article appeared

(b) the statement of permission for use of copyrighted material (see Appendix B: Permissions)

(c) the reprint(s), copied single-sided onto the required type of paper

2. Supplementary materials such as data tables, graphs, and maps which might ordinarily appear as appendices to a thesis.

These two types of appendices form a single sequence, assigned letters and titled as described in this manual. All Appendix pages are part of the single pagination sequence of the thesis. The page numbers will be typed in as needed.

Appendix B

Permissions

Use of copyrighted material in your thesis, including illustrations, usually requires written permission from the copyright holder. Start this time-consuming process as early as possible. Play it safe and assume that you must obtain permission if the material is copyrighted. Consult your advisor or departmental graduate secretary about this process.

Exceptions, sometimes pertaining to small fractions of a musical score or other document, are governed by the concept of ”fair use.” Factors weighed in determining ”fair use” include: the purpose of the use, whether commercial or non-profit and educational; the nature of the copyrighted work; the amount and substance of the material used in relation to the entire work; and the effect of the use upon the potential market for or value of the copyrighted work. The ”fair use” concept is explained in detail in the Chicago Manual of Style. According to the Association of American University Presses, permission is required for quotations which are complete units, for example, an entire poem, letter, book chapter, or an entire map, chart, drawing or other illustration.

Permission to use copyrighted material should be in writing and retained by the author. The release letters should indicate that permission extends to microfilming and publication by University Microfilms Incorporated and that the copyright owners are aware that UMI may sell, on demand, single copies of the thesis, dissertation or document, including the copyrighted materials, for scholarly purposes. UMI requires copies of permission letters to be attached to the publication agreement, and assumes no liability for copyright violations. If permission letters are not supplied, copyrighted materials may not be filmed.

It is polite and good practice to obtain permission to use noncopyrighted material, which may or may not be acknowledged in the text.

For additional information, telephone the Copyright Public Information Office in Washington, DC, (202) 479-0700, weekdays between 8:30 a.m. and 5:00 p.m. EST or write to the Copyright Office, Library of Congress, Washington D.C. 20559.

Appendix C

Human/Animal Subjects Approval

Research involving human subjects or live vertebrate animals requires permission from the relevant University committee. Consult your research director for details. If you are working on a project for which your director has obtained the required permissions, be sure your name is listed on the protocol approval and that you have the control number of the approval in your records.

Research activities involving the use of human subjects require the review and approval of the University Human Subjects Committee. A copy of the Human Subjects approval letter along with The Human Subjects Research Statement must be in the student’s file in the Graduate College Degree Certification Office. Questions regarding protocol can be answered by the Human Subjects Committee. Their telephone number is (602) 626-6721.

Research involving any live vertebrate animals must be approved by the Institutional Animal Care and Use Committee (IACUC) - The Animal Research Protocol Review form must be completed by the student/instructor and submitted to the protocol office for review and approval. Contact University Animal Care for instructions, forms and protocol. Their telephone number is (602) 621-3454

Appendix D

Illustrations, Tables, Graphs

Use illustrative material drawn or computer-generated in black. Color will reproduce in microfilm as shades of grey; use color only if essential to convey a significant point in your work. Material may be laser-printed or drawn in waterproof, permanent ink.

• Use labels or symbols rather than color to identify lines on a graph

• Use cross-hatching rather than color to distinguish areas on a map

• Same margin requirements as Body of Paper

• Place the top of a horizontally-oriented page on the left; the page number should appear in the normal position (the upper right corner of the rotated sheet)

• Printed page numbers are required If the caption is so long that it will not fit on the page with the illustration or table, place it on its own numbered page immediately preceding the page it describes.

Appendix E

Oversized Materials

Reduce oversized pages, such as maps and pictures, to 8.5 by 11 inches without sacrificing legibility. If you must include oversized pages, two options are available:

1. Include a page 11 inches high, but wider than 8.5 inches, folded once or twice. The left edge of the foldout page should be even with the other pages of the thesis, and all folds should be made vertically. The folds on the right must be at least 1 inch from the right edge of other thesis pages to avoid damage to the foldout when pages are trimmed for binding. Place the page number in the upper right corner.

2. If the oversized material cannot be reduced enough for the first option, roll it and submit in a sturdy mailing tube. The document must be labeled in the lower right corner with:

• Name

• Degree

• Department

• Title of thesis

• Date degree awarded

Use a separate tube for each copy and include the label information on the outside of each tube. When you make arrangements with any business to have copies of your thesis bound for departmental or private use, discuss procedures concerning folding of the oversize material and any additional fees which may be incurred for this service. In the LIST OF ILLUSTRATIONS use the phrase, in pocket, instead of a page number.

Appendix F

Photographs

Photographs should be high contrast, low gloss black and white pictures. They may be printed on 8.5 x 11 inch photographic paper in order to avoid mounting.

If mounting of small photographs on a standard page is necessary, use double-sided transparent tape or photographic dry mounting tissue and mount the photographs on #67 pliable bristol board. High quality screen prints or photocopies are acceptable.

References

[1] CCIT, Center for Computing & Information Technology, offers workshops and lectures on word processing, graphs and tables, and thesis/dissertation formatting.

[2] Your advisor and your committee can help; providing assistance on thesis formatting, a part of good mentoring.

[3] Doctoral Candidates: A Handbook For Completing the Steps To Your Degree. Available at no cost from the Graduate College Degree Certification Office.

[4] Master’s and Specialists Candidates: A Handbook For Completing the Steps To Your Degree. Available at no cost from the Graduate College Degree Certification Office.

[5] Publishing Your Dissertation: How to Prepare Your Manuscript for Publication. UMI Dissertation Services. Available at no cost from the Graduate College Degree Certification Office.

[6] Publishing Your Masters Thesis: How to Prepare Your Manuscript for Publication. UMI Dissertation Services. Available at no cost from the Graduate College Degree Certification Office.

[7] Publication style manuals for disciplines are available at the Main Library.